OpenStudy (anonymous):

Let f(x) = 5x + 12. Find f–1(x).

Parth (parthkohli):

1) Find \(x\) in terms of \(y\). \( \color{Black}{\Rightarrow y = 5x + 12}\) \( \color{Black}{\Rightarrow y - 12 = 5x }\) \( \color{Black}{\Rightarrow \Large{y - 12 \over 5} = x}\) 2) Replace \(y\) with \(x\) and \(x\) with \(f^{-1}(x)\)

OpenStudy (anonymous):

y = 5x + 12 x = y - 12 / 5 Ans: x - 12 / 5

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {x - 12 \over 5} = f^{-1}(x) }\)

OpenStudy (jiteshmeghwal9):

ohhhhhhh, so it's\[f^{-1}(x)\]